OpenStudy (anonymous):

2^(x + 1) = 5^(1 - 2x)

OpenStudy (anonymous):

I know that the first step is to put ln on each side of the equation, but i'm not sure where to go after that

OpenStudy (anonymous):

Hit both sides with log

OpenStudy (anonymous):

Aplly log on both sides log 2^(x + 1) = log 5^ ( 1 - 2x)

OpenStudy (anonymous):

log 2^a = a log 2

OpenStudy (anonymous):

so it should be (x + 1) log 2 = (1 - 2x) log 5

OpenStudy (anonymous):

Are ln and log the same thing?

OpenStudy (shamil98):

you can use ln or log, to solve for x, the result will be the same

OpenStudy (anonymous):

what do i do after i have (x + 1) ln or log 2 = (1 - 2x) ln or log 5

OpenStudy (anonymous):

I got (ln 2) x + 1 ln 2 = (-2ln5)x = x - ln 5, but i dont think that's right